# Emanuel Castelo Aix-Marseille Université, CNRS, LIS, 13009 Marseille, France Oscar Defrain Aix-Marseille Université, CNRS, LIS, 13009 Marseille, France Guilherme C. M. Gomes LIRMM, Université de Montpellier, CNRS, Montpellier, France Universidade Federal de Minas Gerais, Belo Horizonte, Brazil

### v2

@@ 1 Introduction
The problem Dom-Enum has gained lots of interest since it has been shown, perhaps surprisingly, to be polynomially equivalent to Trans-Enum, even when restricted to cobipartite graphs [KLMN14]. In particular, this work has initiated a fruitful line of research of characterizing which graph classes admit output-polynomial algorithms, in the hope of better understanding the underlying structure that makes Trans-Enum tractable. For example, polynomial-delay algorithms were obtained for various graph classes including chordal graphs [KLM+15a] (with linear delay on some subclasses [KLMN14, DN19]), graphs of bounded degeneracy [EGM03], line graphs [KLM+15b], graphs of bounded clique-width [Cou09], or graph classes related to permutation graphs and their generalizations [KLMN12, GHK+18, BDMN20]. Incremental-polynomial time algorithms were obtained for graphs of bounded conformality [KBEG07], $C_{\leq 6}$-free graphs [GHKV15], chordal bipartite graphs [GHK+16a], $(C_{4},C_{5},\text{claw})$-free graphs [KLMN12], or graphs arising in geometrical contexts [ERR19, GHK+18] including unit-disks and unit-square graphs. Output-polynomial time algorithms are known for $\log n$-degenerate graphs [EGM03], or graphs of bounded clique number [BDH+20].
@@
Regarding total dominating sets, a first observation is that they are transversals of open neighborhoods. This implies an incremental quasi-polynomial time algorithm for the problem using the results of Fredman and Khachiyan [FK96]. However, few output-polynomial time algorithms are known for specific instances of TDom-Enum. Among them, we cite the class of chordal bipartite graphs, which admits a polynomial-delay algorithm for the problem [GHK+16a]
[…]
Note that, in the literature, particular attention has been put into characterizing the status of these problems on bipartite, split, and cobipartite graphs. This is because such instances contain the incidence graph of any hypergraph, which is fundamentally challenging and makes interesting links with Trans-Enum. However, and in contrast to the split and cobipartite cases, the case of bipartite graphs has been particularly challenging with the question of the (in)tractability of these problems raised in various works [KN14, GHK+16a]. Only recently an output-polynomial time algorithm has been provided for Dom-Enum in a generalization of bipartite graphs [BDH+20]. Still the problem is not known to admit a polynomial-delay algorithm. Concerning TDom-Enum and CDom-Enum, it is also open, to the best of our knowledge, whether they can be solved with polynomial delay, or even in output-polynomial time, in bipartite graphs. Among natural subclasses of bipartite graphs, the case of chordal bipartite graphs has been studied as it admits nice structural properties analogous to chordal graphs, such as perfect elimination orderings and linearly many minimal separators. We note, however, that even in this particular class, the existence of a polynomial-delay algorithm for Dom-Enum seems open, and has been left open in [GHK+16a].
@@
[[GHK+16a]]theoremthmtds The minimal total dominating sets of a chordal bipartite graph be enumerated with polynomial delay.
@@
Concerning CDom-Enum, we show that the minimal separators of a chordal bipartite graph define a hypergraph of bounded conformality (Lemma 5), a parameter introduced by Berge [Ber84] and that has been successfully used in the design of tractable algorithms for domination [KLMN12] and minimal transversals enumeration [KBEG07]. This, combined with the fact that chordal bipartite graphs have polynomially many minimal separators, which can be listed in polynomial time (see Section 2), allow us to obtain an incremental-polynomial time algorithm for CDom-Enum in that class (Theorem 4).
@@ 2 Preliminary notions
We assume the reader to be familiar with standard terminology from graph and hypergraph theory, and refer to the two monographs [Die12, Ber84] for notations not included below.
Graphs. All graphs considered in this paper are simple and finite. If $v$ is a vertex of $G$ then we denote by $N(v)$ its open neighborhood, by $N[v]$ its closed neighborhood, and define its distance-2 neighborhood as $N^{2}(v):=N[N[v]]\setminus N[v]$. The first two notions are naturally extended to subsets by $N[S]:=\bigcup_{v\in S}N[S]$ and $N(S):=(\bigcup_{v\in S}N[S])\setminus S$. An independent set is a set of pairwise non-adjacent vertices. A graph is bipartite if its vertex set $V(G)$ can be partitioned into two independent sets $X,Y$. It is called bipartite chain if, in addition, there exist orderings $x_{1},\dots,x_{p}$ and $y_{1},\dots,y_{q}$ of $X$ and $Y$, respectively, such that $N(x_{i})\subseteq N(x_{j})$ for all $1\leq i\leq j\leq p$, and $N(y_{i})\supseteq N(y_{j})$ for all $1\leq i\leq j\leq q$.
Hypergraphs. Let $\mathcal{H}$ be a hypergraph. Given a vertex $v\in V(\mathcal{H})$, the set of edges that contain $v$, also called edges incident to $v$, is denoted by $\operatorname{{\sf inc}}(v)$. A transversal of $\mathcal{H}$ is a subset $T\subseteq V(\mathcal{H})$ such that $E\cap T\neq\emptyset$ for all $E\in E(\mathcal{H})$, i.e., every edge is incident to (or hit by) a vertex of $T$. The hypergraph induced by a set $X\subseteq V(\mathcal{H})$ of vertices is defines as $\mathcal{H}[X]=(X,\{E\in E(\mathcal{H}):E\subseteq X\})$. It is minimal if it is minimal by inclusion, or equivalently, if $T\setminus\{v\}$ is not a transversal for any $v\in T$. The set of minimal transversals of $\mathcal{H}$ is denoted by $Tr(\mathcal{H})$. The private edges of an element $v$ in a transversal $T$ are those edges of $\mathcal{H}$ which are only hit by $v$ in $T$; we denote the set of such edges by $\operatorname{{\sf priv}}(T,v)$. It is well-known (see e.g., [Ber84]) that a set $T\subseteq V(\mathcal{H})$ is a minimal transversal of $\mathcal{H}$ if and only if $T$ is a transversal and every $v$ in $T$ has a private edge. We formally state the problem of enumerating minimal transversals of a hypergraph as follows.
@@
Note that in the problem above, $\mathcal{H}$ can be considered inclusion-wise minimal (or Sperner) as this does not change its set of minimal transversals [EMG08].
@@
Minimal separators. A separator in a graph $G$ is a subset $S\subseteq V(G)$ such that $G-S$ is disconnected. In the following, let $\mathcal{S}(G):=\{S\subseteq V:G-S~{}\text{is not connected and}~{}\forall S^{\prime}\subset S,G-S^{\prime}~{}\text{is connected}\}$ denote the set of (inclusion-wise) minimal separators of $G$. Then we have the following.
@@
A vertex $v$ in a graph $G$ is called weak-simplicial if $N(v)$ is an independent set, and for any two $x,y\in N(v)$, we have $N(x)\subseteq N(y)$ or $N(y)\subseteq N(x)$; two such vertices $x,y$ are said to be comparable. A weak-simplicial elimination ordering of $G$ is an ordering $v_{1},\dots,v_{n}$ of its vertices such that, for all $1\leq i\leq n$, $v_{i}$ is weak-simplicial in $G_{i}=G\left[\{v_{i},\dots,v_{n}\}\right]$. The following characterization is due to [KWUA19] and may be regarded as a simplification of an earlier characterization and recognition algorithm by Uehara [Ueh02].
@@
Additionally, from the above definition it is easily seen that if $G$ is chordal bipartite, and $v$ is a weak-simplicial vertex, then the graph $G[N^{2}(v)]$ induced by the distance-2 neighborhood of $v$ is bipartite chain; see [KWUA19, Lemma 3]. A characterization based on separators is given in [GG78], where it is proved that a bipartite graph is chordal bipartite if and only if every minimal edge separator induces a complete bipartite subgraph. In [KLP11], the following analogue of the forward implication is given for minimal $a$-$b$ separators (hence for minimal separators).
@@
Finally, we will need two extra properties of minimal separators in general and in chordal bipartite graphs. Let $S$ be a separator of $G$. We say that $C$ is close to $S$ if every vertex in $S$ has at least one neighbor in $C$. It is folklore that minimal $a$-$b$ separators $S$ are characterized by the existence of two close components in $G-S$; see [SL97, Gol04, KLP11] or [KK94] for a proof. We derive one implication for minimal separators since, in particular, they are $a$-$b$ minimal separators for some specific pairs $a,b$ of vertices. Note however that the converse is not true.
@@
Let $G$ be chordal bipartite with bipartition $(A,B)$. Let $S$ be a minimal separator of $G$ and $C$ be a component of $G-S$ that is close to $S$. If $S\cap A\neq\emptyset$ then there exists a vertex $x$ in $C$ with $N(x)\cap S=S\cap A$.
@@ 3 The sequential method
We describe the sequential method. As by Propositions 13, Dom-Enum, TDom-Enum, and CDom-Enum can be seen as particular instances of Trans-Enum we choose to follow the presentation of [BDM24], and refer the reader to [EGM03, EHR08, BDH+20] for more details on this approach.
Let $\mathcal{H}\subseteq 2^{V}$ be a hypergraph on vertex set $V=\{v_{1},\dots,v_{n}\}$. For $1\leq i\leq n$ we denote by $V_{i}:=\{v_{1},\dots,v_{i}\}$ the set of its first $i$ vertices, by $\mathcal{H}_{i}:=\mathcal{H}[V_{i}]$ the hypergraph induced by $V_{i}$, and set $V_{0}:=\mathcal{H}_{0}:=Tr(H_{0}):=\emptyset$. In addition, for $S\subseteq V_{i}$ and $v\in S$, we note $\operatorname{{\sf priv}}_{i}(S,v)$ the set of private edges of $v$ with respect to $S$ in $\mathcal{H}_{i}$. Given a minimal transversal $T\in Tr(\mathcal{H}_{i+1})$ for $i<n$, we denote by $\operatorname{{\sf parent}}(T,i+1)$ the parent of $T$ with respect to $i+1$ obtained after repeatedly removing from $T$ the vertex of smallest index $v$ such that $\operatorname{{\sf priv}}_{i}(T,v)=\emptyset$. By definition, every $T$ has a unique parent. Lastly, we denote by $\operatorname{{\sf children}}(T^{\star},i)$, for $T^{\star}\in Tr(\mathcal{H}_{i})$, the family of $T\in Tr(\mathcal{H}_{i+1})$ such that $T^{\star}=\operatorname{{\sf parent}}(T,i+1)$.
@@
Let $10\leq i\leq n-$ 1 and $T\in Tr(\mathcal{H}_{i+1})$, then $\operatorname{{\sf parent}}(T,i+1)\in Tr(\mathcal{H}_{i})$.
@@
Let $10\leq i\leq n-$ 1 and $T^{\star}\in Tr(\mathcal{H}_{i})$, then either:
@@
Let $f,s:\mathbb{N}^{2}\rightarrow\mathbb{Z}_{+}$ be two functions. If there is an algorithm that, given a hypergraph $\mathcal{H}$ on vertex set $V=\{v_{1},\dots,v_{n}\}$ and $m$ edges, an integer $i\in\{1,\dots,n-1\}$, and $T^{\star}\in Tr(\mathcal{H}_{i})$, enumerates $\operatorname{{\sf children}}(T^{\star},i)$ with delay $f(n,m)$ and using $s(n,m)$ space, then there is an algorithm that enumerates $Tr(\mathcal{H})$ with $\mathcal{O}(nf(n,m)$ delay and $\mathcal{O}(ns(n,m))$ space.
@@
$\Delta_{i+1}:=\{E\in\mathcal{H}_{i+1}:E\cap T^{\star}=\emptyset\}\subseteq\operatorname{{\sf inc}}_{i+1}(v_{i+1})$
@@
As a corollary, in order to enumerate $\operatorname{{\sf children}}(T^{\star},i)$ for some $i\in\{1,\dots,n-1\}$ and $T^{\star}\in Tr(\mathcal{H}_{i})$, one can simply enumerate all the minimal transversals of $\Delta_{i+1}$ and filter out those that do not provide children of $T^{\star}$. In the following, we call minimal extensions of $(T^{\star},i)$ such sets. Note, however that the number of minimal transversal of $\Delta_{i+1}$ may be greater (typically exponential) than the number of actual children, hence that this approach fails to efficiently produce children in general. Moreover, this approach is intractable in general, with instances admitting non-trivial children if and only if NP-complete problems admit a positive answer; see [BDM24] or Section 6. We nevertheless have the following which says that the approach is still tractable if we can bound the number of minimal extension by a polynomial in the input size, and which is in essence what is done in [EGM03].

### v3

@@ 1 Introduction
The problem Dom-Enum has gained lots of interest since it has been shown, perhaps surprisingly, to be polynomially equivalent to Trans-Enum, even when restricted to cobipartite graphs [KLMN14]. In particular, this work has initiated a fruitful line of research of characterizing which graph classes admit output-polynomial algorithms, in the hope of better understanding the underlying structure that makes Trans-Enum tractable. For example, polynomial-delay algorithms were obtained for various graph classes including chordal graphs [KLM+15a] (with linear delay on some subclasses [KLMN14, DN19]), graphs of bounded degeneracy [EGM03], line graphs [KLM+15b], graphs of bounded clique-width [Cou09], or graph classes related to permutation graphs and their generalizations [KLMN12, GHK+18, BDMN20]. Incremental-polynomial time algorithms were obtained for graphs of bounded conformality [KBEG07], $C_{\leq 6}$-free graphs [GHKV15], chordal bipartite graphs [GHK+16a], $(C_{4},C_{5},\text{claw})$-free graphs [KLMN12], or graphs arising in geometrical contexts [ERR19, GHK+18] including unit-disk and unit-square graphs. Output-polynomial time algorithms are known for $\log n$-degenerate graphs [EGM03], and graphs of bounded clique number [BDH+20].
@@
Regarding total dominating sets, a first observation is that they are transversals of open neighborhoods. This implies an incremental quasi-polynomial time algorithm for the problem using the results of Fredman and Khachiyan [FK96]. However, few output-polynomial time algorithms are known for specific instances of TDom-Enum. Among them, we cite the class of chordal bipartite graphs, which admits a polynomial-delay algorithm for the problem [GHK+16a].
Concerning connected dominating sets, it is not known whether the problem reduces to Trans-Enum. It has in fact been conjectured that this is not the case in [BBHK15]. Rather, it has been shown in [KLMN14] that the problem reduces to the enumeration of minimal transversals of an implicit hypergraph given by the minimal separators of the graph. However, the number of such separators may be exponential, and are intractable to compute [BDK+24]. The only work known to us dealing with connected dominating sets enumeration, apart from [KLMN14], falls under the input-sensitive approach [GHK16b, GHKS20, AKFG+22] where one aims at small exponential times. Still, we note that some essential questions concerning output-sensitive enumeration were addressed in [AKFG+22] such as the $\mathsf{NP}$-hardness of the associated extension problem, i.e., deciding if a given set $U\subseteq V(G)$ is contained in some minimal connected dominating set.
Note that, in the literature, particular attention has been put into characterizing the status of these problems on bipartite, split, and cobipartite graphs. This is because such instances contain the incidence graph of any hypergraph, which is fundamentally challenging and makes interesting links with Trans-Enum. However, and in contrast to the split and cobipartite cases, the case of bipartite graphs has been particularly challenging with the question of the (in)tractability of these problems raised in various works [KN14, GHK+16a]. Only recently has an output-polynomial time algorithm been provided for Dom-Enum in a generalization of bipartite graphs [BDH+20]. Still the problem is not known to admit a polynomial-delay algorithm. Concerning TDom-Enum and CDom-Enum, it is also open, to the best of our knowledge, whether they can be solved with polynomial delay, or even in output-polynomial time, in bipartite graphs. Among natural subclasses of bipartite graphs, the case of chordal bipartite graphs has been studied as it admits nice structural properties analogous to chordal graphs, such as perfect elimination orderings and linearly many minimal separators. We note, however, that even in this particular class, the existence of a polynomial-delay algorithm for Dom-Enum seems open, and has been left open in [GHK+16a].
@@
[[GHK+16a]]theoremthmtds The minimal total dominating sets of a chordal bipartite graph can be enumerated with polynomial delay.
@@
Concerning CDom-Enum, we show that the minimal separators of a chordal bipartite graph define a hypergraph of bounded conformality (Lemma 5), a parameter introduced by Berge [Ber84] and successfully used in the design of tractable algorithms for domination [KLMN12] and minimal transversals enumeration [KBEG07]. This, combined with the fact that chordal bipartite graphs have polynomially many minimal separators, which can be listed in polynomial time (see Section 2), allows us to obtain an incremental-polynomial time algorithm for CDom-Enum in that class (Theorem 4).
@@ 2 Preliminary notions
We assume familiarity with standard terminology from graph and hypergraph theory, and refer to [Die12, Ber84] for notations not included below.
Graphs. All graphs considered in this paper are simple and finite. Given $X\subseteq V(G)$, $G[X]\coloneq(X,\{uv\in E(G)\mid u,v\in X\})$ is the subgraph of $G$ induced by $X$. Given $v\in V(G)$, we denote by $N(v)$ its open neighborhood, by $N[v]$ its closed neighborhood, and define its distance-2 neighborhood as $N^{2}(v)\coloneq N[N[v]]\setminus N[v]$. The first two notions are naturally extended to subsets by $N[S]\coloneq\bigcup_{v\in S}N[S]$ and $N(S)\coloneq(\bigcup_{v\in S}N[S])\setminus S$. An independent set is a set of pairwise non-adjacent vertices. A graph is bipartite if its vertex set $V(G)$ can be partitioned into two independent sets $X,Y$. It is called bipartite chain if, in addition, there exist orderings $x_{1},\dots,x_{p}$ and $y_{1},\dots,y_{q}$ of $X$ and $Y$, respectively, such that $N(x_{i})\subseteq N(x_{j})$ for all $1\leq i\leq j\leq p$, and $N(y_{i})\supseteq N(y_{j})$ for all $1\leq i\leq j\leq q$. Note that bipartite chain graphs are closed under taking induced subgraphs.
Hypergraphs. Let $\mathcal{H}$ be a hypergraph. Given a vertex $v\in V(\mathcal{H})$, the set of edges that contain $v$, also called edges incident to $v$, is denoted by $\operatorname{{\sf inc}}(v)$. The hypergraph induced by a set $X\subseteq V(\mathcal{H})$ of vertices is defined as $\mathcal{H}[X]=(X,\{E\in E(\mathcal{H}):E\subseteq X\})$. A transversal of $\mathcal{H}$ is a subset $T\subseteq V(\mathcal{H})$ such that $E\cap T\neq\emptyset$ for all $E\in E(\mathcal{H})$, i.e., every edge is incident to (or hit by) a vertex of $T$. It is minimal if it is minimal by inclusion, or equivalently, if $T\setminus\{v\}$ is not a transversal for any $v\in T$. The set of minimal transversals of $\mathcal{H}$ is denoted by $Tr(\mathcal{H})$. The private edges of an element $v$ in a transversal $T$ are those edges of $\mathcal{H}$ which are only hit by $v$ in $T$; we denote the set of such edges by $\operatorname{{\sf priv}}(T,v)$. It is well-known (see e.g., [Ber84]) that a set $T\subseteq V(\mathcal{H})$ is a minimal transversal of $\mathcal{H}$ if and only if $T$ is a transversal and every $v$ in $T$ has a private edge. We formally state the problem of enumerating minimal transversals of a hypergraph as follows.
@@
Note that in the problem above, $\mathcal{H}$ can be considered inclusion-wise minimal (or Sperner) as this does not change its set of minimal transversals, and removing non-minimal hyperedges can be performed in polynomial time [EMG08].
@@
Minimal separators. A separator in a graph $G$ is a subset $S\subseteq V(G)$ such that $G-S$ is disconnected. In the following, let $\mathcal{S}(G)\coloneq\{S\subseteq V:G-S~\text{is not connected and}~\forall S^{\prime}\subset S,G-S^{\prime}~\text{is connected}\}$ denote the set of (inclusion-wise) minimal separators of $G$. Then we have the following.
@@
A vertex $v$ in a graph $G$ is called weak-simplicial if $N(v)$ is an independent set, and for any two $x,y\in N(v)$, we have $N(x)\subseteq N(y)$ or $N(y)\subseteq N(x)$; two such vertices $x,y$ are said to be comparable. Note that this is precisely to say that the graph $G[N(v)\cup N^{2}(v)]$ induced by the vertices at distance at most 222 from $v$ is a bipartite chain; see [KWUA19, Lemma 3]. A weak-simplicial elimination ordering of $G$ is an ordering $v_{1},\dots,v_{n}$ of its vertices such that, for all $1\leq i\leq n$, $v_{i}$ is weak-simplicial in $G_{i}=G\left[\{v_{1},\dots,v_{i}\}\right]$. The following characterization is due to [KWUA19] and may be regarded as a simplification of an earlier characterization and recognition algorithm by Uehara [Ueh02].
@@
Another characterization of chordal bipartite graphs based on separators is given in [GG78], where it is proved that a bipartite graph is chordal bipartite if and only if every minimal edge separator induces a complete bipartite subgraph. In [KLP11], the following analogue of the forward implication is given for minimal $a$-$b$ separators (hence for minimal separators).
@@
Finally, we will need two extra properties of minimal separators in general and in chordal bipartite graphs. Let $S$ be a separator of $G$ and $C$ be a component of $G-S$. We say that $C$ is close to $S$ if every vertex in $S$ has at least one neighbor in $C$. It is folklore that minimal $a$-$b$ separators $S$ are characterized by the existence of two close components in $G-S$; see [SL97, Gol04, KLP11] or [KK94] for a proof. We derive one implication for minimal separators since, in particular, they are $a$-$b$ minimal separators for some specific pairs $a,b$ of vertices. Note however that the converse is not true.
@@
[…]
@@ 3 The sequential method
We describe the sequential method. As by Propositions 13, the problems Dom-Enum, TDom-Enum, and CDom-Enum can be seen as particular instances of Trans-Enum we choose to follow the presentation of [BDM24], and refer the reader to [EGM03, EHR08, BDH+20] for more details on this approach.
Let $\mathcal{H}\subseteq 2^{V}$ be a hypergraph on vertex set $V=\{v_{1},\dots,v_{n}\}$. For $1\leq i\leq n$ we denote by $V_{i}\coloneq\{v_{1},\dots,v_{i}\}$ the set of its first $i$ vertices, by $\mathcal{H}_{i}\coloneq\mathcal{H}[V_{i}]$ the hypergraph induced by $V_{i}$, and set $V_{0}\coloneq\mathcal{H}_{0}\coloneq Tr(\mathcal{H}_{0})\coloneq\emptyset$. In addition, for $S\subseteq V_{i}$ and $v\in S$, we note $\operatorname{{\sf priv}}_{i}(S,v)$ the set of private edges of $v$ with respect to $S$ in $\mathcal{H}_{i}$. Given a minimal transversal $T\in Tr(\mathcal{H}_{i+1})$ for $i<n$, we denote by $\operatorname{{\sf parent}}(T,i+1)$ the parent of $T$ with respect to $i+1$ obtained after repeatedly removing from $T$ the vertex of smallest index $v$ such that $\operatorname{{\sf priv}}_{i}(T,v)=\emptyset$. By definition, every $T$ has a unique parent. Lastly, we denote by $\operatorname{{\sf children}}(T^{\star},i)$, for $T^{\star}\in Tr(\mathcal{H}_{i})$, the family of $T\in Tr(\mathcal{H}_{i+1})$ such that $T^{\star}=\operatorname{{\sf parent}}(T,i+1)$.
@@
Let $0\leq i\leq n-1$ and $T\in Tr(\mathcal{H}_{i+1})$, then $\operatorname{{\sf parent}}(T,i+1)\in Tr(\mathcal{H}_{i})$.
@@
Let $0\leq i\leq n-1$ and $T^{\star}\in Tr(\mathcal{H}_{i})$, then either:
@@
Let $f,s:\mathbb{N}^{2}\rightarrow\mathbb{Z}_{+}$ be two functions. If there is an algorithm that, given a hypergraph $\mathcal{H}$ on vertex set $V=\{v_{1},\dots,v_{n}\}$ and $m$ edges, an integer $i\in\{1,\dots,n-1\}$, and $T^{\star}\in Tr(\mathcal{H}_{i})$, enumerates $\operatorname{{\sf children}}(T^{\star},i)$ with delay $f(n,m)$ and using $s(n,m)$ space, then there is an algorithm that enumerates $Tr(\mathcal{H})$ with $O(nf(n,m))$ delay and $O(ns(n,m))$ space.
@@
$\Delta_{i+1}\coloneq\{E\in\mathcal{H}_{i+1}:E\cap T^{\star}=\emptyset\}\subseteq\operatorname{{\sf inc}}_{i+1}(v_{i+1})$
@@
As a corollary, in order to enumerate $\operatorname{{\sf children}}(T^{\star},i)$ for some $i\in\{1,\dots,n-1\}$ and $T^{\star}\in Tr(\mathcal{H}_{i})$, one can simply enumerate all the minimal transversals of $\Delta_{i+1}$ and filter out those that do not provide children of $T^{\star}$. In the following, we call such sets minimal extensions of $(T^{\star},i)$. Note, however that the number of minimal transversal of $\Delta_{i+1}$ may be greater (typically exponential) than the number of actual children, hence that this approach fails to efficiently produce children in general. Moreover, this approach is intractable in general, with instances admitting non-trivial children if and only if NP-complete problems admit a positive answer; see [BDM24] or Section 6. We nevertheless have the following which says that the approach is still tractable if we can bound the number of minimal extension by a polynomial in the input size, and which is in essence what is done in [EGM03].